# How Can Time Series Analysis Benefit From Multiple Modalities? A Survey and Outlook

### v2

@@ 1. Introduction
Recently, multiple ”richer” modalities such as language and vision have recently experienced explosive development, not only as separate fields but also in multi-modal study, represented by powerful foundation multi-modal models, such as GPT and Qwen series. In contrast, the time-series modality remains relatively underexplored and isolated. We notice that many recent TSA studies have begun breaking this impasse (Ming et al., 2023; Liu et al., 2024f, c, 2025b; Merrill et al., 2024; Wang et al., 2024b; Xie et al., 2024b; Kong et al., 2025), driven by a unified high-level motivation: ”How can TSA benefit from multiple modalities?”
@@
Our survey serves as the first review of the emerging MM4TSA field and systematically identifies three key approaches: (1) TimeAsX: Reusing Foundation Models from Other Modalities for Efficient TSA; (2) Time+X: Multimodal Extensions for Enhanced TSA; and (3) Time2X and X2Time: Cross-Modality Interaction for Advanced TSA. Additionally, we comprehensively cover multiple modalities, including language, vision, tables, and audio, by grouping existing works according to modality type. Moreover, we introduce representative studies from specific domains, i.e., finance, medical, and spatiotemporal, to clearly illustrate the real-world value of these three approaches. Through this taxonomy, we initially identify key gaps associated with each approach: (1) which modality to reuse, (2) how to handle heterogeneous modality combinations, and (3) how to generalize to unseen tasks. We discuss potential solutions to inspire future work. Additionally, we compile a list of representative datasets and maintain an up-to-date GitHub repository of the discussed papers to facilitate future research.
@@
SFT (Xie et al., 2024b); System 1 vs 2 (Liu et al., 2025b) , fill=gray!5 ] ] ] [Domain Specific (§5.5), fill=red!20 [Spatial-Temporal Data, fill=red!10 [(Wang et al., 2017; Yang et al., 2019; Irvin et al., 2024; Yan et al., 2024; Li et al., 2024b), fill=purple!0] ] [Medical TS, fill=red!10 [(Bleich et al., 2024; Wan et al., 2024b; Tang et al., 2024b; Wan et al., 2024a; Li et al., 2024a; Hunter et al., 2008b; Gatt et al., 2009; Hunter et al., 2008a)
(Alcaraz and Strodthoff, 2023; Chung et al., 2023; Lai et al., 2025; Zhao et al., 2024b; Guo et al., 2024; Yu et al., 2024; Liu et al., 2023c; Chan et al., 2024; Cosentino et al., 2024), fill=purple!0] ] [Financial TS, fill=red!10 [(Liu et al., 2022a, 2023a; Chen et al., 2024b; Li et al., 2023a; Bhatia et al., 2024; Xie et al., 2024c), fill=purple!0] ] ] [ Other Cross-Modality Works (§5.4, fill=green!20 [(Islam et al., 2024) ,fill=purple!0] ] [Text 2 Time + Time 2 Text (§5.3), fill=green!20 [(Wang et al., 2024b; Xie et al., 2024b; Kong et al., 2025; Kim et al., 2024b) ,fill=purple!0] ] [Text 2 Time (§5.1), fill=green!20 [Generation, fill=orange!10 [(Lai et al., 2025; Merrill et al., 2024; Xie et al., 2024b), fill=purple!0] ] [Retrieval, fill=orange!10 [(Fons et al., 2024; Cai et al., 2024; Ito et al., 2024), fill=purple!0] ] ] [Time 2 Text (§5.2), fill=green!20 [Explanation, fill=orange!10 [(Aksu et al., 2024; Jiang et al., 2025; Liu et al., 2024i), fill=purple!0] ] [Captioning, fill=orange!10 [(Sharma et al., 2021; Spreafico and Carenini, 2020; Li et al., 2023b; Zhang et al., 2023b; Dohi et al., 2024; Lin et al., 2024; Trabelsi et al., 2025), fill=purple!0] ] ] ] [Time+X:
@@
Robust (Ebrahimi et al., 2023) , fill=gray!5 ] ] ] [Domain Specific (§4.3), fill=red!20 [Spatial-Temporal Data, fill=red!10 [(Anantharam et al., 2016; Fan et al., 2022; Lin et al., 2022; Chen et al., 2024a; Feng et al., 2024; Lin et al., 2023; Zhang et al., 2024d; Han et al., 2024; Jiang et al., 2024; Kosugi et al., 2022), fill=purple!0] ] [Medical TS, fill=red!10 [(Wagner et al., 2020; Lee et al., 2023; Wang et al., 2024c; Zhu et al., 2024; Yang and Wu, 2021; Niu et al., 2023; King et al., 2023; Huang et al., 2020; Kyono et al., 2021)
@@
(Liu et al., 2024d; Darapaneni et al., 2022; Tavakoli et al., 2025; Rekabsaz et al., 2017; Bamford et al., 2023; Dong et al., 2024), fill=purple!0] ] ] [Time + Other Modalities (§4.2), fill=green!20 [(Ebrahimi et al., 2023; Girdhar et al., 2023), fill=purple!0] ] [Time + Text (§4.1), fill=green!20 [(Ming et al., 2023; Liu et al., 2024j; Williams et al., 2024; Wu et al., [n. d.]; Wang et al., 2024b; Liu et al., 2025a; Xu et al., 2024a; Xinlei et al., 2024; Cheng et al., 2024; Tao et al., 2024; Liu et al., 2024f; Li et al., 2025b; Zhang et al., 2024c), fill=purple!0] ] ] [TimeAsX:
@@
Composite (Ruan et al., 2025) , fill=gray!5 ] ] ] [Domain Specific (§3.4), fill=red!20 [Spatial-Temporal Data, fill=red!10 [(Liu et al., 2024h; Zhang et al., 2024b; Liu et al., 2024g; Yuan et al., 2024; Li et al., 2024b, 2025c; Yao et al., 2018)
@@
(Liu et al., 2024i; Xie et al., 2024b; Hu et al., 2025; Pan et al., 2024; Ansari et al., 2024; Wang et al., 2024b; Chen et al., 2025; Huang et al., 2025) , fill=purple!0] ] ] ]
@@ 2. Background and Taxonomy
A time series, an ordered sequence of data points, can be denoted as $X_{1:T}=\{x_{1},x_{2},\dots,x_{T}\}$. Classic TSA tasks include forecasting, which predicts future values based on historical observations (e.g., flu forecasting (Mathis et al., 2024)); classification, which assigns time series into discrete categories (e.g., ECG signal classification (Kiranyaz et al., 2015)); anomaly detection, which identifies abnormal patterns (e.g., financial fraud detection (Ren et al., 2019)); and imputation, which addresses missing data (e.g., completing medical records (Bertsimas et al., 2021)). In this survey, we exclude generalized time series such as videos (Oprea et al., 2020) or event streams (Li et al., 2024c), since these data are inherently multimodal. More details about background are provided in Appendix A.
@@ 3.1. Time As Text
Given the success of large language models (LLMs) in a wide range of language understanding and generation tasks, recent works have explored the usage of LLMs for TSA tasks (Gruver et al., 2023; Liu et al., 2024j; Xinglei et al., 2023; Hao et al., 2022; Zhou et al., 2023; Rasul et al., 2023; Yuxuan et al., 2024; Ching et al., 2023; Chenxi et al., 2023; Cao et al., 2024; Ming et al., 2023; Liu et al., 2024e, i; Xie et al., 2024b; Hu et al., 2025; Pan et al., 2024; Ansari et al., 2024; Wang et al., 2024b; Chen et al., 2025; Huang et al., 2025). The motivation behind these works is mainly based on the fact that both language and time series have a sequence structure, and on the belief that the general ability of LLMs can be applied to TSA.
@@ 3.1.3. Training for Alignment with Expanded Vocabulary
Some methods expand LLM vocabulary to align with time-series datasets: they treat time-series data as sentences in a foreign language and adapt the LLMs toward such a language. These works differ in how they design the adaptor functions to map time-series to expanded vocabulary. Chronos (Ansari et al., 2024) quantizes the normalized input time-series values into discrete tokens that are fed into a language model. Wang et al. (2024b) similarly introduce additional tokens for quantized values of input time-series as well as NaN or missing values.
@@ 3.4. Domain-Specific Time-Series Works
A majority of spatial-temporal works aim to transform time series into textual prompts and thus introduce large language models (LLMs) as a powerful forecaster (Liu et al., 2024h; Zhang et al., 2024b; Liu et al., 2024g; Yuan et al., 2024; Li et al., 2024b). For example, UrbanGPT (Li et al., 2024b) proposes to convert time series to textual sequences and fine-tune an LLM with instructions containing decomposed spatial-temporal descriptions plus brief task descriptions. Details of works within this line are provided in Appendix F. Besides, some also propose to convert spatial-temporal data to images (Yao et al., 2018; Li et al., 2022; Bao et al., 2022; Jiang et al., 2023; Tang et al., 2024a). Specifically, they view the input data as a sequence of images corresponding to each time frame, then use vision models to encode each image to obtain stronger spatial understanding. Details of works within this line are provided in Appendix G.
@@
ChatTime (Wang et al., 2024b) Time+Text 3 Real Datasets (weather&date)
@@ 4. Time+X: Multimodal Extension for
Human experts typically complete TSA tasks by integrating multiple modalities, especially by combining numerical and textual data. For instance, epidemiologists combine influenza infections data with textual domain knowledge, policies, and reports to predict future epidemiological trends. However, most TSA models (Vaswani et al., 2017; Liu et al., 2023b; Zhang and Yan, 2022; Liu et al., 2022b; Zhou et al., 2022; Kitaev et al., 2020; Zhou et al., 2021; Nie et al., 2022; Liu et al., 2024b) are unimodal, using only numerical series. Extending unimodal TSA to multimodal TSA, especially through the integration of text, is an emerging topic (Ming et al., 2023; Liu et al., 2024j; Wei et al., 2022; Williams et al., 2024; Jin et al., 2023; Wu et al., [n. d.]; Wang et al., 2024b; Liu et al., 2025a; Xu et al., 2024a; Dong et al., 2024; Liu et al., 2024b; Xinlei et al., 2024; Cheng et al., 2024; Tao et al., 2024; Liu et al., 2024f; Li et al., 2025b; Zhang et al., 2024c). We first introduce multimodal TSA for general domains, starting with Time+Text and then Time+Other Modalities, followed by specific domains, using representative works. We summarize some representative multimodal time series datasets in Table 1.
@@ 4.1.1. Integrated Text Types
A direct way to add exogenous information is to include dates, such as weekends or holidays, or weather details, such as precipitation or sunrise time (Wang et al., 2024b; Xu et al., 2024a; Liu et al., 2025a). This type of information is easy to get. For example, Wang et al. pairs each time-series sample with corresponding text such as: ”…The target date for prediction is Monday, December 5, 2022. It is a weekday with moderate snow fall and light breeze…”
@@ 4.3. Domain-Specific Time-Series Works
In the medical domain, time-series analysis is often combined with text, image, and tabular data. The most widely used datasets are MIMIC (Time-Series + Text + Image + Tabular) and PTB-XL (Time-Series + Text). MIMIC-Based Datasets, with MIMIC-IV (Johnson et al., 2020) and MIMIC-III (Johnson et al., 2016), containing electronic health records (EHRs), physiological time-series (vitals, lab results), and free-text clinical notes, supporting tasks like mortality prediction, ICU event forecasting, and disease progression. PTB-XL (Wagner et al., 2020) provides electrocardiogram (ECG) waveforms annotated with medical reports, enabling multimodal research that pairs time-series ECG signals with textual diagnosis descriptions for clinical classification tasks. Image integration is widely studied, particularly in radiology, where physiological signals like electrocardiograms (ECGs) and patient vital signs are analyzed with medical images like X-rays, magnetic resonance imaging (MRI), and computed tomography (CT) scans to enhance diagnostic accuracy and clinical predictions (Wang et al., 2024c; Yao et al., 2024; Hayat et al., 2022). Time-series and tabular data fusion focuses on structured EHR data, combining static demographic, genetic, or laboratory test results with dynamic patient monitoring signals for predictive modeling (Soenksen et al., 2022; Zhang et al., 2023a; Kyung et al., 2024). Each modality combination serves specific medical tasks as detailed in Appendix J.0.1
@@ 5.3. Time 2 Text + Text 2 Time
These works focus on multimodal multitask time series models, specifically handling text+time inputs and outputs, characterized by general time-series question & answering (QA). ChatTime (Wang et al., 2024b) reinterprets time series as a foreign language by normalizing, discretizing, and expanding the vocabulary of a pre-trained language model, thereby enabling bimodal inputs and outputs for multiple TSA tasks including zero-shot unimodal or multimodal forecasting and time-series QA. ChatTS Xie et al. bridges the gap between time series data and natural language by aligning LLMs with synthetic data. ChatTS is also designed for multimodal multivariate input and output for general-purpose time series tasks. Kong et al. propose Time-MQA to enable comprehensive time series analysis by unifying diverse tasks—including forecasting, imputation, anomaly detection, classification, and open-ended reasoning—within a single natural language question answering framework. Time-MQA is designed to accept text+time input and produce text (time-series included) outputs for general-purpose TSA tasks across various domains. Kim et al. focus on the joint text and time series forecasting tasks and introduce the Hybrid Multi-Modal Forecaster.

### v3

@@ Abstract.
Figure 1. Illustration of three approaches of multiple modalities for TSA (MM4TSA). These three MM4TSA approaches empower TSA from modality reusing, multimodal enhancement, and cross-modal interaction, respectively.
@@ 1. Introduction
Recently, multiple ”richer” modalities such as language and vision have recently experienced explosive development, not only as separate fields but also in multi-modal study, represented by powerful foundation multi-modal models, such as GPT and Qwen series. In contrast, the time-series modality remains relatively underexplored and isolated. We notice that many recent TSA studies have begun breaking this impasse (Ming et al., 2023; Liu et al., 2024f, c, 2025b; Merrill et al., 2024; Wang et al., 2024c; Xie et al., 2024b; Kong et al., 2025), driven by a unified high-level motivation: ”How can TSA benefit from multiple modalities?”
@@
Our survey serves as the first review of the emerging MM4TSA field and systematically identifies three key approaches, as illustrated in Figure 1: (1) TimeAsX: Reusing Foundation Models from Other Modalities for Efficient TSA; (2) Time+X: Multimodal Extensions for Enhanced TSA; and (3) Time2X and X2Time: Cross-Modality Interaction for Advanced TSA. Additionally, we comprehensively cover multiple modalities, including language, vision, tables, and audio, by grouping existing works according to modality type. Moreover, we introduce representative studies from specific domains, i.e., finance, medical, and spatiotemporal, to clearly illustrate the real-world value of these three approaches. Through this taxonomy, we initially identify key gaps associated with each approach: (1) which modality to reuse, (2) how to handle heterogeneous modality combinations, and (3) how to generalize to unseen tasks. We discuss potential solutions to inspire future work. Additionally, we compile a list of representative datasets and maintain an up-to-date GitHub repository of the discussed papers to facilitate future research.
@@
SFT (Xie et al., 2024b); System 1 vs 2 (Liu et al., 2025b) , fill=gray!5 ] ] ] [Domain Specific (§5.5), fill=red!20 [Spatial-Temporal Data, fill=red!10 [(Wang et al., 2017; Yang et al., 2019; Irvin et al., 2024; Yan et al., 2024; Li et al., 2024c), fill=purple!0] ] [Medical TS, fill=red!10 [(Bleich et al., 2024; Wan et al., 2024b; Tang et al., 2024b; Wan et al., 2024a; Li et al., 2024b; Hunter et al., 2008b; Gatt et al., 2009; Hunter et al., 2008a)
(Alcaraz and Strodthoff, 2023; Chung et al., 2023; Lai et al., 2025; Zhao et al., 2024b; Guo et al., 2024; Yu et al., 2024; Liu et al., 2023c; Chan et al., 2024; Cosentino et al., 2024), fill=purple!0] ] [Financial TS, fill=red!10 [(Liu et al., 2022a, 2023a; Chen et al., 2024b; Li et al., 2023a; Bhatia et al., 2024; Xie et al., 2024c), fill=purple!0] ] ] [ Other Cross-Modality Works (§5.4, fill=green!20 [(Islam et al., 2024) ,fill=purple!0] ] [Text 2 Time + Time 2 Text (§5.3), fill=green!20 [(Wang et al., 2024c; Xie et al., 2024b; Kong et al., 2025; Kim et al., 2024b) ,fill=purple!0] ] [Text 2 Time (§5.1), fill=green!20 [Generation, fill=orange!10 [(Lai et al., 2025; Merrill et al., 2024; Xie et al., 2024b), fill=purple!0] ] [Retrieval, fill=orange!10 [(Fons et al., 2024; Cai et al., 2024; Ito et al., 2024), fill=purple!0] ] ] [Time 2 Text (§5.2), fill=green!20 [Explanation, fill=orange!10 [(Aksu et al., 2024; Jiang et al., 2025; Liu et al., 2024i), fill=purple!0] ] [Captioning, fill=orange!10 [(Sharma et al., 2021; Spreafico and Carenini, 2020; Li et al., 2023b; Zhang et al., 2023b; Dohi et al., 2024; Lin et al., 2024; Trabelsi et al., 2025), fill=purple!0] ] ] ] [Time+X:
@@
Robust (Ebrahimi et al., 2023) , fill=gray!5 ] ] ] [Domain Specific (§4.3), fill=red!20 [Spatial-Temporal Data, fill=red!10 [(Anantharam et al., 2016; Fan et al., 2022; Lin et al., 2022; Chen et al., 2024a; Feng et al., 2024; Lin et al., 2023; Zhang et al., 2024d; Han et al., 2024; Jiang et al., 2024; Kosugi et al., 2022), fill=purple!0] ] [Medical TS, fill=red!10 [(Wagner et al., 2020; Lee et al., 2023; Wang et al., 2024d; Zhu et al., 2024; Yang and Wu, 2021; Niu et al., 2023; King et al., 2023; Huang et al., 2020; Kyono et al., 2021)
@@
(Liu et al., 2024d; Darapaneni et al., 2022; Tavakoli et al., 2025; Rekabsaz et al., 2017; Bamford et al., 2023; Dong et al., 2024), fill=purple!0] ] ] [Time + Other Modalities (§4.2), fill=green!20 [(Ebrahimi et al., 2023; Girdhar et al., 2023), fill=purple!0] ] [Time + Text (§4.1), fill=green!20 [(Ming et al., 2023; Liu et al., 2024j; Williams et al., 2024; Wu et al., [n. d.]; Wang et al., 2024c; Liu et al., 2025a; Xu et al., 2024a; Xinlei et al., 2024; Cheng et al., 2024; Tao et al., 2024; Liu et al., 2024f; Li et al., 2025b; Zhang et al., 2024c), fill=purple!0] ] ] [TimeAsX:
@@
Composite (Ruan et al., 2025) , fill=gray!5 ] ] ] [Domain Specific (§3.4), fill=red!20 [Spatial-Temporal Data, fill=red!10 [(Liu et al., 2024h; Zhang et al., 2024b; Liu et al., 2024g; Yuan et al., 2024; Li et al., 2024c, 2025c; Yao et al., 2018)
@@
(Liu et al., 2024i; Xie et al., 2024b; Hu et al., 2025; Pan et al., 2024; Ansari et al., 2024; Wang et al., 2024c; Chen et al., 2025; Huang et al., 2025) , fill=purple!0] ] ] ]
@@ 2. Background and Taxonomy
A time series, an ordered sequence of data points, can be denoted as $X_{1:T}=\{x_{1},x_{2},\dots,x_{T}\}$. Classic TSA tasks include forecasting, which predicts future values based on historical observations (e.g., flu forecasting (Mathis et al., 2024)); classification, which assigns time series into discrete categories (e.g., ECG signal classification (Kiranyaz et al., 2015)); anomaly detection, which identifies abnormal patterns (e.g., financial fraud detection (Ren et al., 2019)); and imputation, which addresses missing data (e.g., completing medical records (Bertsimas et al., 2021)). In this survey, we exclude generalized time series such as videos (Oprea et al., 2020) or event streams (Li et al., 2024d), since these data are inherently multimodal. More details about background are provided in Appendix A.
@@ 3.1. Time As Text
Given the success of large language models (LLMs) in a wide range of language understanding and generation tasks, recent works have explored the usage of LLMs for TSA tasks (Gruver et al., 2023; Liu et al., 2024j; Xinglei et al., 2023; Hao et al., 2022; Zhou et al., 2023; Rasul et al., 2023; Yuxuan et al., 2024; Ching et al., 2023; Chenxi et al., 2023; Cao et al., 2024; Ming et al., 2023; Liu et al., 2024e, i; Xie et al., 2024b; Hu et al., 2025; Pan et al., 2024; Ansari et al., 2024; Wang et al., 2024c; Chen et al., 2025; Huang et al., 2025). The motivation behind these works is mainly based on the fact that both language and time series have a sequence structure, and on the belief that the general ability of LLMs can be applied to TSA.
@@ 3.1.3. Training for Alignment with Expanded Vocabulary
Some methods expand LLM vocabulary to align with time-series datasets: they treat time-series data as sentences in a foreign language and adapt the LLMs toward such a language. These works differ in how they design the adaptor functions to map time-series to expanded vocabulary. Chronos (Ansari et al., 2024) quantizes the normalized input time-series values into discrete tokens that are fed into a language model. Wang et al. (2024c) similarly introduce additional tokens for quantized values of input time-series as well as NaN or missing values.
@@ 3.4. Domain-Specific Time-Series Works
A majority of spatial-temporal works aim to transform time series into textual prompts and thus introduce large language models (LLMs) as a powerful forecaster (Liu et al., 2024h; Zhang et al., 2024b; Liu et al., 2024g; Yuan et al., 2024; Li et al., 2024c). For example, UrbanGPT (Li et al., 2024c) proposes to convert time series to textual sequences and fine-tune an LLM with instructions containing decomposed spatial-temporal descriptions plus brief task descriptions. Details of works within this line are provided in Appendix F. Besides, some also propose to convert spatial-temporal data to images (Yao et al., 2018; Li et al., 2022; Bao et al., 2022; Jiang et al., 2023; Tang et al., 2024a). Specifically, they view the input data as a sequence of images corresponding to each time frame, then use vision models to encode each image to obtain stronger spatial understanding. Details of works within this line are provided in Appendix G.
@@
ChatTime (Wang et al., 2024c) Time+Text 3 Real Datasets (weather&date)
@@ 4. Time+X: Multimodal Extension for
Human experts typically complete TSA tasks by integrating multiple modalities, especially by combining numerical and textual data. For instance, epidemiologists combine influenza infections data with textual domain knowledge, policies, and reports to predict future epidemiological trends. However, most TSA models (Vaswani et al., 2017; Liu et al., 2023b; Zhang and Yan, 2022; Liu et al., 2022b; Zhou et al., 2022; Kitaev et al., 2020; Zhou et al., 2021; Nie et al., 2022; Liu et al., 2024b) are unimodal, using only numerical series. Extending unimodal TSA to multimodal TSA, especially through the integration of text, is an emerging topic (Ming et al., 2023; Liu et al., 2024j; Wei et al., 2022; Williams et al., 2024; Jin et al., 2023; Wu et al., [n. d.]; Wang et al., 2024c; Liu et al., 2025a; Xu et al., 2024a; Dong et al., 2024; Liu et al., 2024b; Xinlei et al., 2024; Cheng et al., 2024; Tao et al., 2024; Liu et al., 2024f; Li et al., 2025b; Zhang et al., 2024c). We first introduce multimodal TSA for general domains, starting with Time+Text and then Time+Other Modalities, followed by specific domains, using representative works. We summarize some representative multimodal time series datasets in Table 1.
@@ 4.1.1. Integrated Text Types
A direct way to add exogenous information is to include dates, such as weekends or holidays, or weather details, such as precipitation or sunrise time (Wang et al., 2024c; Xu et al., 2024a; Liu et al., 2025a). This type of information is easy to get. For example, Wang et al. pairs each time-series sample with corresponding text such as: ”…The target date for prediction is Monday, December 5, 2022. It is a weekday with moderate snow fall and light breeze…”
@@ 4.3. Domain-Specific Time-Series Works
In the medical domain, time-series analysis is often combined with text, image, and tabular data. The most widely used datasets are MIMIC (Time-Series + Text + Image + Tabular) and PTB-XL (Time-Series + Text). MIMIC-Based Datasets, with MIMIC-IV (Johnson et al., 2020) and MIMIC-III (Johnson et al., 2016), containing electronic health records (EHRs), physiological time-series (vitals, lab results), and free-text clinical notes, supporting tasks like mortality prediction, ICU event forecasting, and disease progression. PTB-XL (Wagner et al., 2020) provides electrocardiogram (ECG) waveforms annotated with medical reports, enabling multimodal research that pairs time-series ECG signals with textual diagnosis descriptions for clinical classification tasks. Image integration is widely studied, particularly in radiology, where physiological signals like electrocardiograms (ECGs) and patient vital signs are analyzed with medical images like X-rays, magnetic resonance imaging (MRI), and computed tomography (CT) scans to enhance diagnostic accuracy and clinical predictions (Wang et al., 2024d; Yao et al., 2024; Hayat et al., 2022). Time-series and tabular data fusion focuses on structured EHR data, combining static demographic, genetic, or laboratory test results with dynamic patient monitoring signals for predictive modeling (Soenksen et al., 2022; Zhang et al., 2023a; Kyung et al., 2024). Each modality combination serves specific medical tasks as detailed in Appendix J.0.1. Besides, a large number of multimodal studies have emerged on sequence structure data such as proteins (Xiao et al., 2025), RNA (Xiao et al., 2024) and neuro signal . But these works are beyond the scope of this survey because these data do not have timestamps.
@@ 5.3. Time 2 Text + Text 2 Time
These works focus on multimodal multitask time series models, specifically handling text+time inputs and outputs, characterized by general time-series question & answering (QA). ChatTime (Wang et al., 2024c) reinterprets time series as a foreign language by normalizing, discretizing, and expanding the vocabulary of a pre-trained language model, thereby enabling bimodal inputs and outputs for multiple TSA tasks including zero-shot unimodal or multimodal forecasting and time-series QA. ChatTS Xie et al. bridges the gap between time series data and natural language by aligning LLMs with synthetic data. ChatTS is also designed for multimodal multivariate input and output for general-purpose time series tasks. Kong et al. propose Time-MQA to enable comprehensive time series analysis by unifying diverse tasks—including forecasting, imputation, anomaly detection, classification, and open-ended reasoning—within a single natural language question answering framework. Time-MQA is designed to accept text+time input and produce text (time-series included) outputs for general-purpose TSA tasks across various domains. Kim et al. focus on the joint text and time series forecasting tasks and introduce the Hybrid Multi-Modal Forecaster.